If the product of a number and -7 is reduced by 5, the result is 75 less than twice the opposite of the number. What is the number?
Answer by erica65404(394)

product of a number and -7
-7n
reduced by 5
-7n-5
equals
twice the opposite of the number
-n is the opposite of n
2(-n) or -2n
75 less than -2n
-2n-75
-7n-5=-2n-75
solve for n
-7n=-2n-70
-5n=-70
n=14
The number is 14.